OpenStudy (anonymous):

Why is it important to know how to multiply fractions in order to divide fractions? Give a real-world example when you would need to know how to multiply fractions or mixed numbers in order to solve a problem.

OpenStudy (anonymous):

well if you have to multiply 1/2 cups of flower by 1/3 when cooking you would need to know how to multiply fractions.

OpenStudy (anonymous):

Um that doesnt really explain it

OpenStudy (anonymous):

Tapha help plz

OpenStudy (irishboy123):

**Why is it important to know how to multiply fractions in order to divide fractions? ** because to divide fractions, you turn the second fraction upside down and then multiply eg \(\frac{3}{2} \color{red}{\div} \frac{5}{7} = \frac{3}{2} \color{Red}{\times} \frac{7}{5}\)